OpenStudy (anonymous):

*Will fan and medal!* A bus travels 480 miles in 7 hours and 30 minutes. What was the average rate of travel in miles per hour?

OpenStudy (kkutie7):

you can start here... \[\frac{30min}{1}\frac{1hour}{60min}=?hours\] add it to 7hours

OpenStudy (anonymous):

R=D/T

OpenStudy (kkutie7):

^ yes that is good. once you have the total amount of hours you need to put it in that form. average rate= total distance/time

OpenStudy (anonymous):

The total time of the trip was 5 hours and 45 minutes. ... starts at the beginning of a runners' path and runs at a rate of 6 miles per hour. ... One hour later a bus traveling at 50 MPH leaves for the same town and ... My total distance is 480 miles.

OpenStudy (anonymous):

^^^^^ @queenleviosa